Seitz Glassware is trying to determine its growth rate for an annual cash dividend. The most recent​ dividend, Div 0​, was $ 0.30 per share. The​ stock's target return rate is 12​%. What is the​ stock's price if a.  the annual growth rate is 2​%? b.  the annual growth rate is 5​%? c.  the annual growth rate is 6​%? d.  the annual growth rate is 8​%? e.  the annual growth rate is 11​%?
Answer #1

Using DDM Model,

a.

Annual Growth Rate = 2%

Stock Price = 0.30(1.02)/(0.12 - 0.02)

Stock Price = $3.06

b.

Annual Growth Rate = 6%

Stock Price = 0.30(1.06)/(0.12 - 0.06)

Stock Price = $5.3

c.

Annual Growth Rate = 8%

Stock Price = 0.30(1.08)/(0.12 - 0.08)

Stock Price = $8.1

d.

Annual Growth Rate = 11%

Stock Price = 0.30(1.11)/(0.12 - 0.11)

Stock Price = $33.3